OpenStudy (anonymous):

Assume that y varies inversely as x. If y = -6 when x = -2, find y when x = 5.

OpenStudy (anonymous):

@amistre64 @thomaster

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

an inverse varaince is of the form: y = k/x therefore the constant is: k= yx = -6(-2) = 12 this produces the equation: y = 12/x to play with

OpenStudy (anonymous):

y = 3 ?

OpenStudy (anonymous):

y = 2.4

OpenStudy (amistre64):

i would say yes, then 12/5 = 2.4

OpenStudy (anonymous):

yayy so i got it right!!

OpenStudy (anonymous):

i have one more problem im stuck on if you could help me?

OpenStudy (anonymous):

Use the Law of the Lever to solve the following. Mark sits 4.5 feet from the center of the seesaw to balance a friend who weighs 53 lbs and is sitting 6 feet from the center. How many pounds does Mark weigh? Round to the nearest hundredth.

OpenStudy (amistre64):

law of the lever .. is that like: whoever has the biggest stick wins?

OpenStudy (anonymous):

i have no idea, im terrible at word problems

OpenStudy (amistre64):

Fd = Fd

OpenStudy (anonymous):

im assuming

OpenStudy (amistre64):